I need help solving for variables in a word problem. I am not very good at word problems could someone please walk me through this problem.
The perimeter of a quadrilateral (4 sided polygon) is 29 inches. THe longest side is twice as long as the shortest side. The other two sides are equally long and are 2 inches longer than the shortest side. Find the length of all four sides.
Im not even sure how to pull an equation from this. please help
thank you

Answer by nerdybill(7384)

The idea is to assign a variable to things you don't know.
In this case, every side is related to the "shortest side". So...
.
Let x = length of the shortest side
2x = length of longest side
x+2 = length of the last two sides
.
The problem gives us the perimeter as 29 inches:
x + 2x + 2(x+2) = 29
Now, we solve for x:
x + 2x + 2x + 4 = 29
5x + 4 = 29
5x = 25
x = 5 inches (shortest side)
2x = 2(5) = 10 inches (longest side)
x+2 = 5+2 = 7 inches (the two remaining sides)